The area of a triangle is 18a4b7c square centimeters. What are the possible dimensions for the base and height of the triangle?
Answer by Alan3354(69443)

The area of a triangle is 18a4b7c square centimeters. What are the possible dimensions for the base and height of the triangle?
-----------
Area = b*h/2 = 18a^4b^7c --- I think that's what you mean.
Use ^ for exponents.
------------
b*h = 36a^4b^7c
-----
b = 1, h = 36a^4b^7c
b = 2, h = 18a^4b^7c
b = pi, h = 36a^4b^7c/pi
There are an infinite # of possibilities.